OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (anonymous):

really need help

OpenStudy (anonymous):

-2x + y = 1 i.e. y= 2x+1-----(1 -4x + y = -1------(2 substituting y= 2x+1 in eq (2 we find: -4x + 2x+1 = -1 -2x+1=-1 i.e. -2x= -2 i.e. x= -2/-2 i.e. x= 1 hence y= 2x+1 = 2*1 + 1 = 2+1 = 3 So (x, y) = (1, 3) is required solution set. Hence last option is correct answer. @alyssa_fields50

OpenStudy (anonymous):

@alyssa_fields50

OpenStudy (anonymous):

I have more questions if anyone else like to help me itd be very appreciated

OpenStudy (anonymous):

Go ahead, but medal first

OpenStudy (anonymous):

Solve the system of equations. x + 3y = -1 2x + 2y = 6 (-4, 1) (2, -1) (4, -1) (5, -2) sorry for taking so long my connection keeps going out.